Circular No. 7093

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN M31 R. Stagni, Astronomy Department, University of Padua, reports the

discovery of an apparent nova in M31 by U. Buonomo and F. Di Mille with the 1.20-m reflector of the Asiago Astrophysical Observatory. The nova was found on V and R CCD images taken on Jan. 14.8, 15.8, 16.8 (V ∼ 17.8, R ∼ 17.6), and 18.9 UT. The object is located at α = 0^{h}42^{m}43^{s}.8, δ = +41^{o}11′32′′ (equinox 2000.0). V and R CCD images of M31 taken on 1998 Dec. 14.8 and 17.9 show nothing at the position of the nova (limiting mag V and R ∼ 19.3).

SUPERNOVA 1998fc IN ANONYMOUS GALAXY A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise

Observatory Optical Transients Search (see IAUC 6917, 7055) that a spectrum (range 333–755 nm) of SN 1998fc (IAUC 7082, 7091), obtained by P. Leisy, O. Hainaut, and T. Sekiguchi with the European Southern Observatory (ESO) 3.6-m Telescope (+ EFOSC spectrograph) on Jan. 14, confirms this to be a type-Ia supernova, a few weeks past maximum brightness. The estimated supernova redshift is 0.10, consistent with its being in the galaxy cluster Abell 403 (z = 0.103).

V1333 AQUILAE C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence,

write: “We searched CCD frames, taken in the quiescent state of V1333 Aql = Aql X-1 on 1989 May 8 with the ESO 3.6-m telescope (+ EFOSC; projected pixel size 0′′.33), for the line-of-sight interloper (cf. IAUC 7086). Three consecutive dithered 5-min V -band exposures (seeing 1′′.2) have been analyzed using DAOPHOT psf-fitting routines. A clean psf profile was derived from five nearby isolated stars after three neighbor-removing iterations. Subtraction of this profile as fitted to the image at the location of V1333 Aql reveals a residual faint object located 2 ± 0.5 pixels (0′′.66 ± 0′′.17) to the west of the main object and having the average stellar FWHM for the frame; it appears on each individual frame, and its location is consistent with the 0′′.46 separation between the two images reported on IAUC 7086. None of the other stars of comparable brightness shows a similar residue. We derive V = 22.9 ± 0.3 for the faint object and V = 19.25 ± 0.01 for V1333 Aql. If this residual object is indeed the reported interloper, then the contamination appears to be slight in the V band (3-percent level) and suggests a very late-type spectrum for the interloper.”

1999 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green